# Comparison of Extended Lubrication Theories for Stokes Flow

### v1

@@ 2.3 Extended Lubrication Theory
An alternative extended lubrication theory model (T.G.-ELT) is proposed by Takeuchi and Gu [14], which retains terms of order up to $\varepsilon^{2}$ in Eqs. 4 and 5 to result in the governing equations,
(22) $\frac{\partial p}{\partial x}$ $=\eta\frac{\partial^{2}u}{\partial y^{2}},$
(23) $\frac{\partial p}{\partial y}$ $=\eta\frac{\partial^{2}v}{\partial y^{2}},$
with incompressibility Eq. 3 and the boundary conditions Eqs. 10, 8 and 9. Note that the second-order term $\varepsilon^{2}\frac{\partial^{2}\bar{u}}{\partial\bar{x}^{2}}$ appearing in Eq. 4 is not included in the T.G.-ELT formulation, since they argue this may be neglected in the case of small surface gradients. As with the perturbed solutions, the flux $\mathcal{Q}$ is assumed to be equal the Reynolds flux, and in general, the pressure drop $\Delta\mathcal{P}$ may deviate from that of the solution to the Reynolds equation.
According to the T.G.-ELT assumptions, the pressure gradient $\frac{\partial^{2}p}{\partial y\partial x}$ is negligible provided that the condition $\varepsilon^{2}\ll 1$ is satisfied [14]. Since the $x$ momentum equations of lubrication theory and T.G.-ELT are the same, integration in $y$ of Eq. 22 gives $u=u_{\text{Re}}$ as in Eq. 11, and applying incompressibility gives $v=v_{\text{Re}}$ as in Eq. 12. However, the $y$ momentum equations are different; integration in $y$ of Eq. 23 gives the pressure decomposition,
(24) $p(x,y)=p_{\text{Re}}(x)+p_{\text{Adj}}(x,y)+\eta\varsigma(x),$
where $p_{\text{Adj}}(x,y)=\eta\frac{\partial v_{\text{Re}}}{\partial y}$ and $\varsigma(x)$ is an integral constant. That is,
(25) $p_{\text{Adj}}(x,y)=\frac{-1}{2}\frac{d^{2}p_{\text{Re}}}{dx^{2}}y^{2}+\Bigg(\frac{1}{2}\bigg(\frac{d^{2}p_{\text{Re}}}{dx^{2}}h(x)+\frac{dp_{\text{Re}}}{dx}\frac{dh}{dx}\bigg)-\frac{\mathcal{U\eta}}{\big[h(x)\big]^{2}}\frac{dh}{dx}\Bigg)y.$
@@ 3 Velocity-Adjusted Extended Lubrication Theory
In [14], the integral constant $\varsigma(x)$ is assumed negligible on account of $\varepsilon^{2}\ll 1$. In the construction of the velocity to follow, we remove this assumption and allow $\varsigma(x)\neq 0$ to ensure the velocity derived from $p(x,y)$ satisfies the boundary conditions and incompressibility, even in regions where the T.G.-ELT conditions are violated. The inclusion of the term $\varsigma(x)$ leads to $p$, $u$, and $v$ each having different expressions compared with the T.G.-ELT solution; we refer to the new solution with $\varsigma(x)\neq 0$ as the velocity adjusted extended lubrication theory solution (VA-ELT).
@@
Since $p_{\text{Adj}}(x,y)$ is polynomial in $y$, the $x$ momentum equation Eq. 22 may be integrated exactly in $y$ to obtain $u$. We utilize the integral constant $\varsigma(x)$ to impose the boundary conditions and incompressibility independent of the local geometry. Applying the boundary conditions Eq. 8 for $u$,
(26) $u(x,y)=u_{\text{Re}}(x,y)-\frac{1}{24\eta}\frac{d^{3}p_{\text{Re}}}{dx^{3}}(y^{4}-h^{3}y)+\frac{1}{12\eta}\frac{d^{2}}{dx^{2}}\Bigg[h\frac{dp_{\text{Re}}}{dx}\Bigg](y^{3}-h^{2}y)\\ +\frac{\mathcal{U}}{6}\frac{d^{2}}{dx^{2}}\Bigg[\frac{1}{h}\Bigg](y^{3}-h^{2}y)+\frac{1}{2}\frac{d\varsigma}{dx}(y^{2}-hy).$
@@
Then applying incompressibility and integrating in $y$ with the boundary conditions Eq. 9 for $v$,
(27) $v(x,y)=v_{\text{Re}}(x,y)+\frac{1}{24\eta}\Bigg(\frac{d^{4}p_{\text{Re}}}{dx^{4}}\bigg(\frac{y^{5}}{5}-\frac{h^{3}y^{2}}{2}\bigg)-\frac{3h^{2}y^{2}}{2}\frac{d^{3}p_{\text{Re}}}{dx^{3}}\frac{dh}{dx}\Bigg)\\ -\frac{1}{12\eta}\Bigg(\frac{d^{3}}{dx^{3}}\bigg[h\frac{dp}{dx}\bigg]\bigg(\frac{y^{4}}{4}-\frac{h^{2}y^{2}}{2}\bigg)-\frac{d^{2}}{dx^{2}}\bigg[h\frac{dp_{\text{Re}}}{dx}\bigg]\frac{dh}{dx}hy^{2}\Bigg)\\ +\frac{\mathcal{U}}{6}\Bigg(\frac{y^{4}-2h^{2}y^{2}}{4h^{2}}\frac{d^{3}h}{dx^{3}}+\frac{3y^{4}-2h^{2}y^{2}}{2h^{4}}\bigg[\frac{dh}{dx}\bigg]^{3}+\frac{4h^{2}y^{2}-3y^{4}}{2h^{3}}\frac{dh}{dx}\frac{d^{2}h}{dx^{2}}\Bigg)\\ -\frac{d^{2}\varsigma}{dx^{2}}\bigg(\frac{y^{3}}{6}-\frac{hy^{2}}{4}\bigg)+\frac{d\varsigma}{dx}\frac{dh}{dx}\frac{y^{2}}{4}.$
@@
Observe that if we were to assume $\varsigma(x)=0$, the velocity Eqs. 26 and 27 is not necessarily incompressible and Eq. 27 may not satisfy $v(x,h)=0$.
The integral constant $\varsigma(x)$ is determined by asserting a constant flux. We prescribe the flux $\mathcal{Q}=\mathcal{Q}_{\text{Re}}$ as in Eq. 13. Then $\varsigma(x)$ satisfies,
(28) $\frac{d\varsigma}{dx}=\frac{3h^{2}}{20\eta}\frac{d^{3}p_{\text{Re}}}{dx^{3}}-\frac{h}{4\eta}\bigg(h\frac{d^{3}p_{\text{Re}}}{dx^{3}}+2\frac{d^{2}p_{\text{Re}}}{dx^{2}}\frac{dh}{dx}+\frac{dp_{\text{Re}}}{dx}\frac{d^{2}h}{dx^{2}}\bigg)+\frac{\mathcal{U}}{2}\bigg(\frac{1}{h}\frac{d^{2}h}{dx^{2}}-\frac{2}{h^{2}}\bigg[\frac{dh}{dx}\bigg]^{2}\Bigg).$
@@
For an arbitrary height function, numerical integration is required to obtain $\varsigma(x)$. The boundary condition $\varsigma(x_{L})=0$ corresponds to $p(x_{L},y)=p_{\text{Re}}(x_{L},y)=0$. The total pressure drop is then $\Delta\mathcal{P}=\Delta\mathcal{P}_{\text{Re}}+\varsigma(x_{0})$.
@@ 4 Results
We now compare the lubrication models for various geometries and examine the solutions of velocity and pressure for a variety of textured slider examples. The Reynolds, T.G.-ELT, VA-ELT, $\varepsilon^{2}$-PLT and $\varepsilon^{4}$-PLT solutions are compared against the Stokes solution Appendix B. The accuracy of pressure and velocity solutions are characterized using the relative percent error in $L_{2}$ norm. For an appropriate comparison of solutions from the various models, the flux $\mathcal{Q}$ is kept constant and the pressure drops $\Delta\mathcal{P}$ are presumed to vary.
Note, for the T.G.-ELT method, the prescribed flux $\mathcal{Q}$ corresponds to the pressure gradient $\frac{\partial p}{\partial x}$ assigned at the inlet. However – subject to the magnitude of surface variation – this flux may not be conserved throughout the domain because incompressibility is not strictly enforced.
@@ 4.1 The logistic step
where $H_{\text{in}}\geq H_{\text{out}}>0$ and $\lambda>0$. The logistic step allows us to vary the magnitude of surface gradient while keeping a fixed $\varepsilon=L_{y}/L_{x}$ and step expansion ratio $H_{\text{in}}/H_{\text{out}}$. We consider solutions for $2\leq\lambda\leq 32$, and with fixed $H_{\text{in}}=2$, $H_{\text{out}}=1$, $L=16$, $\nu=1$, $\mathcal{U}=0$ and $\mathcal{Q}=1$. Then the length scale ratio is $\varepsilon\simeq 1/8$, and the maximum magnitude surface gradient is $\big|\frac{dh}{dx}\big|=\lambda/4$.
@@
Figure 2: Pressure and velocity solutions for the logistic step with $\lambda=8$. The ELT and $\varepsilon^{k}$-PLT solutions overestimate $\frac{\partial p}{\partial y}$ in the vicinity of the surface variation, leading to flow recirculation which is not observed in the Stokes solution at this moderate $\lambda$.
The errors in velocity and pressure compared with the Stokes solution are shown in Fig. 3. For the wide range of $\lambda<32$, the VA-ELT solution of pressure is significantly more accurate than the Reynolds and T.G.-ELT solutions (Fig. 3). The T.G.-ELT solution of pressure yields a marginal improvement over the Reynolds pressure. For $2<\lambda<8$, the $\varepsilon^{2}$-PLT solution is best at approximating the pressure, and the VA-ELT solution is best at approximating the velocity. The $\varepsilon^{2}$-PLT solution performs better than the $\varepsilon^{4}$-PLT solution for all $\lambda>2$. For $\lambda\geq 32$, the Reynolds solution has the smallest error in velocity and pressure; the Reynolds solution has the most stable error across the range of $\lambda$. Overall, both the $\varepsilon^{2}$-PLT and VA-ELT solutions perform well for the logistic step up to moderate surface variation. We note however, that all the lubrication models discussed here break down in the limit $\lambda\to\infty$ of a discontinuous step, see Appendix D.
@@ 4.2 The triangular slider
(30) $h(x)=\begin{cases}H_{\text{in}}&x\leq L_{\text{in}}\\ H-\frac{H-H_{\text{in}}}{L_{\text{a}}}\big(L_{\text{in}}+L_{a}-x\big)&L_{\text{in}}\leq x\leq L_{\text{in}}+L_{a}\\ H_{\text{out}}-\frac{H_{\text{out}}-H}{L_{b}}\big(L-L_{\text{out}}-x\big)&L_{\text{in}}+L_{a}\leq x\leq L-L_{\text{out}}\\ H_{\text{out}}&L-L_{\text{out}}\leq x\leq L.\end{cases}$
We consider varying central vertex heights $\frac{1}{16}\leq H\leq 2$, and with fixed $L_{a}=1.25$, $L_{b}=0.75$, $L_{\text{in}}=L_{\text{out}}=7$, $H_{\text{in}}=H_{\text{out}}=1$, $\nu=1$, $\mathcal{U}=0$ and $\mathcal{Q}=1$. The maximum magnitude surface gradient is $|\frac{dh}{dx}|=|1-H|/0.75$, and the length scale ratio is $\varepsilon\simeq\max(1,H)/16$. This configuration is similarly considered in [15] for the case of negative texturing $(H<1)$ with the $\varepsilon^{k}$-PLT solutions.
@@
The pressure and velocity solutions to the triangular slider with $H=0.5$ (following [15]) and $H=2$ are shown in Figs. 5 and 6 respectively. This class of examples demonstrates the limitations of the lubrication models in the case of a discontinuous surface gradient. At the discontinuities in $\frac{dh}{dx}$, the pressure gradient $\frac{d^{2}p_{\text{Re}}}{dx^{2}}$ and velocity $v_{\text{Re}}$ of the Reynolds solution are not necessarily continuous. Consequently, the pressure and velocity solutions from ELT and $\varepsilon^{k}$-PLT may not be continuous at the discontinuities of $\frac{dh}{dx}$. Note that $u_{\text{Re}}$ and $\frac{dp_{\text{Re}}}{dx}$ are continuous for the triangular slider because $h(x)$ is continuous; owing to the lubrication assumptions, $v_{\text{Re}}$ is typically small enough in magnitude compared to $u_{\text{Re}}$ that the discontinuity in $v_{\text{Re}}$ is not significant. However, for larger $H$, the discontinuity in $v_{\text{Re}}$ is more pronounced and leads to a noticeable loss of smoothness in the ELT and $\varepsilon^{k}$-PLT solutions.
In the case of a sufficiently small interior angle at the triangle apex, the Stokes solution is known to exhibit a sequence of eddies receding into the corner [11, 3]. For $H=2$ as in Fig. 6, the first eddy is clearly observed in the Stokes solution. For the ELT and $\varepsilon^{k}$-PLT solutions, flow recirculation may occur if $H$ and $\frac{dh}{dx}$ are sufficiently large. However, because the velocities for these models are not continuous at the discontinuities of $\frac{dh}{dx}$, corner eddies do not occur as in the Stokes solution. In fact, as seen in Fig. 6 with the $\varepsilon^{k}$-PLT solutions, the velocity in the steeper half of the triangle exhibits flow recirculation at smaller $H$ than the shallower half. That is, if $H$ is large enough and $\frac{dh}{dx}$ is discontinuous in magnitude, the Reynolds, ELT and $\varepsilon^{k}$-PLT solutions may exhibit spontaneous flow reversal at the discontinuities of $\frac{dh}{dx}$.
(a) Pressure contours
(b) Velocity streamlines
@@
Figure 5: Pressure and velocity solutions for the triangular slider with $H=0.5$. The Reynolds, ELT and PLT models overestimate the maximum velocity. Discontinuities in velocity and pressure of the ELT and $\varepsilon^{k}$-PLT models are less pronounced at smaller $H$.
(a) Pressure contours
(b) Velocity streamlines
@@
Figure 6: Pressure and velocity solutions for the triangular slider with $H=2$. The velocity and pressure solutions of ELT and $\varepsilon^{k}$-PLT are discontinuous at the discontinuities of $\frac{dh}{dx}$. The sequence of corner eddies characteristic of the Stokes solution is not observed in any of the lubrication models.
@@
The errors in pressure and velocity compared with the Stokes solution for the triangular slider are shown in Fig. 7. The error in velocity is very similar for all of the models. For pressure, the VA-ELT solution is a consistent improvement compared to the Reynolds and T.G.-ELT solutions, and the $\varepsilon^{k}$-PLT solutions perform better still. For negative texturing ($H<1$), the $\varepsilon^{k}$-PLT solutions of pressure have significantly smaller error relative to both the VA-ELT and Reynolds solutions; for positive texturing ($H>1$), the $\varepsilon^{k}$-PLT and VA-ELT solutions are more similar. The error in both pressure and velocity decreases as the difference between the minimum and maximum height decreases. In particular, for the case of negative texturing, the increasing magnitude of surface gradient as the minimum height decreases corresponds to an increase in error for pressure and velocity; this emphasizes the importance of considering the magnitude of surface gradients as well as the length scale ratio when selecting a fluid model.
@@
Figure 7: Relative percent error in velocity $(u,v)$ and pressure $p(x,y)$ compared with the Stokes solution for the triangular slider with varying apex height $H$. Decreased difference between maximum and minimum height corresponds to decreased error in pressure and velocity.

### v2

@@ 2.3 Extended Lubrication Theory
An alternative extended lubrication theory model (T.G.-ELT) is proposed by Takeuchi & Gu [takeuchi_extended_2019], which retains terms of order up to $\varepsilon^{2}$ in Eqs. 4 and 5 to result in the governing equations,
(23) $\frac{\partial P}{\partial X}$ $=\frac{\partial^{2}U}{\partial Y^{2}},$
(24) $\frac{\partial P}{\partial Y}$ $=\varepsilon^{2}\frac{\partial^{2}V}{\partial Y^{2}},$
with incompressibility Eq. 3. Note that the second-order term $\varepsilon^{2}\frac{\partial^{2}U}{\partial X^{2}}$ appearing in Eq. 4 is not included in the T.G.-ELT formulation, since they argue this term may be neglected in the case of small surface gradients [takeuchi_extended_2019]. Moreover, the pressure gradient $\frac{\partial^{2}P}{\partial Y\partial X}$ is considered negligible provided $\varepsilon^{2}\ll 1$ [takeuchi_extended_2019].
Since the $X$ momentum equations of lubrication theory and T.G.-ELT are the same and $\frac{\partial^{2}P}{\partial Y\partial X}$ is considered negligible, integration of Eq. 23 with the boundary conditions Eq. 6 gives $U=U_{0}$ as in Eq. 12. Then applying incompressibility and integrating with the boundary conditions Eq. 7 gives $V=V_{0}$ as in Eq. 13. However, the $Y$ momentum equations are different; integration of Eq. 24 gives the pressure decomposition,
(25) $P(X,Y)=P_{0}(X)+P_{\text{Adj}}(X,Y)+\varsigma(X),$
where $P_{\text{Adj}}(x,y)=\varepsilon^{2}\frac{\partial V_{0}}{\partial Y}$ and $\varsigma(X)$ is an integral constant. Differentiating $V_{0}$ from Eq. 13 yields,
(26) $P_{\text{Adj}}(x,y)=\varepsilon^{2}\Bigg(\frac{-1}{2}\frac{d^{2}P_{0}}{dX^{2}}Y^{2}+\Bigg(\frac{1}{2}\bigg(\frac{d^{2}P_{0}}{dX^{2}}H+\frac{dP_{0}}{dX}\frac{dH}{dX}\bigg)-\frac{U_{b}}{H^{2}}\frac{dH}{dX}\Bigg)Y\Bigg).$
We assume the boundary condition Eq. 15 for $P_{0}$, and $P(1,0)=0$. As with the PLT solution, note that the higher-order approximation $P$ does not necessarily satisfy the pressure gradient condition in Eq. 15. Because the velocity is assumed to satisfy $U\equiv U_{0}$, the flux $\mathcal{Q}$ is the same as that of the order $\varepsilon^{0}$ solution. The average pressure drop is determined by the solution $P$, and in general $\Delta P_{\text{T.G.-ELT}}$ will differ from $\Delta P_{0}$. In [takeuchi_extended_2019], the integral constant $\varsigma$ is assumed negligible on account of $\varepsilon^{2}\ll 1$ and the assumption of small surface gradients. For $\varsigma=0$, the average pressure drop is given by $\Delta P_{\text{ T.G.-ELT}}=\Delta P_{0}+\Delta P_{\text{Adj}}$. Since $P_{\text{Adj}}$ is polynomial in $Y$, $\Delta P_{\text{Adj}}$ may be evaluated exactly from Eq. 9 for a given $H$.
@@ 3 Velocity-Adjusted Extended Lubrication Theory
In the approach of Takeuchi & Gu [takeuchi_extended_2019], the integral constant $\varsigma$ is assumed negligible. In the construction of the velocity to follow, we replace this assumption by requiring $\varsigma$ be chosen to ensure the velocity satisfies the boundary conditions and incompressibility, regardless of whether the T.G.-ELT conditions are satisfied. Including the potentially non-zero term $\varsigma$ leads to additional terms in $P$, $U$, and $V$ compared with the T.G.-ELT solution. We refer to the new solution including $\varsigma$ as the velocity adjusted extended lubrication theory solution (VA-ELT).
@@
With $P$ of the form Eq. 25 and $P_{\text{Adj}}$ given by Eq. 26, the $X$ momentum equation Eq. 23 may be integrated exactly in $Y$ to obtain $U$. We utilize the integral constant $\varsigma$ to impose the boundary conditions and incompressibility independent of the local geometry. Applying the boundary conditions Eq. 6 for $U$ gives,
(27) $U(X,Y)=\varepsilon^{2}\Bigg(\frac{-1}{24}\frac{d^{3}P_{0}}{dX^{3}}(Y^{4}-H^{3}Y)+\frac{1}{12}\frac{d^{2}}{dX^{2}}\bigg[H\frac{dP_{0}}{dX}\bigg](Y^{3}-H^{2}Y)\\ +\frac{U_{b}}{6}\frac{d^{2}}{dX^{2}}\bigg[\frac{1}{H}\bigg](Y^{3}-H^{2}Y)\Bigg)+\frac{1}{2}\frac{d\varsigma}{dX}(Y^{2}-HY)+U_{0}(X,Y).$
@@
Then applying incompressibility Eq. 3 to $U$ in Eq. 27, and integrating with the boundary conditions Eq. 7 for $V$ gives,
(28) $V(X,Y)=\varepsilon^{2}\Bigg(\frac{1}{24}\bigg(\frac{d^{4}P_{0}}{dX^{4}}\bigg(\frac{Y^{5}}{5}-\frac{H^{3}Y^{2}}{2}\bigg)-\frac{d^{3}P_{0}}{dX^{3}}\frac{dH}{dX}\frac{3H^{2}Y^{2}}{2}\bigg)\\ -\frac{1}{12}\bigg(\frac{d^{3}}{dX^{3}}\bigg[H\frac{dP_{0}}{dX}\bigg]\bigg(\frac{Y^{4}}{4}-\frac{H^{2}Y^{2}}{2}\bigg)-\frac{d^{2}}{dX^{2}}\bigg[H\frac{dP_{0}}{dX}\bigg]\frac{dH}{dX}HY^{2}\bigg)\\ +\frac{U_{b}}{6}\bigg(\frac{d^{3}H}{dX^{3}}\frac{Y^{4}-2H^{2}Y^{2}}{4H^{2}}+\bigg[\frac{dH}{dX}\bigg]^{3}\frac{3Y^{4}-2H^{2}Y^{2}}{2H^{4}}+\frac{dH}{dX}\frac{d^{2}H}{dX^{2}}\frac{4H^{2}Y^{2}-3Y^{4}}{2H^{3}}\bigg)\Bigg)\\ -\frac{d^{2}\varsigma}{dX^{2}}\bigg(\frac{Y^{3}}{6}-\frac{HY^{2}}{4}\bigg)+\frac{d\varsigma}{dX}\frac{dH}{dX}\frac{Y^{2}}{4}+V_{0}(X,Y).$
@@
Observe that if we assume $\varsigma=0$, the velocity field from Eqs. 27 and 28 is not necessarily incompressible, and Eq. 28 may not satisfy $V(X,H)=0$.
The pressure $P_{0}$ is assumed to satisfy the boundary conditions Eq. 15. For the higher-order pressure, we assume $P(1,0)=0$, but note that $P$ does not necessarily satisfy the pressure gradient condition in Eq. 15. The integral constant $\varsigma$ is determined by requiring a constant flux, and as a consequence the average pressure drop is determined by Eq. 9. We prescribe $\mathcal{Q}$ to be consistent with the flux of the order $\varepsilon^{0}$ solution. Then $\varsigma$ satisfies,
(29) $\frac{d\varsigma}{dX}=\varepsilon^{2}\Bigg(\frac{3H^{2}}{20}\frac{d^{3}P_{0}}{dX^{3}}-\frac{H}{4}\bigg(H\frac{d^{3}P_{0}}{dX^{3}}+2\frac{d^{2}P_{0}}{dX^{2}}\frac{dH}{dX}+\frac{dP_{0}}{dX}\frac{d^{2}H}{dX^{2}}\bigg)\\ +\frac{U_{b}}{2}\bigg(\frac{1}{H}\frac{d^{2}H}{dX^{2}}-\frac{2}{H^{2}}\bigg[\frac{dH}{dX}\bigg]^{2}\bigg)\Bigg).$
@@
For an arbitrary height function, numerical integration is required to obtain $\varsigma$. The boundary condition $P(1,0)=0$ corresponds to $\varsigma(1)=0$. The average pressure drop is then $\Delta P_{\text{ VA-ELT}}=\Delta P_{0}+\Delta P_{\text{Adj}}+\varsigma(0)$.
@@ 4 Results
We now compare the lubrication models for various geometries and examine the solutions of velocity and pressure for a variety of textured slider examples. The Reynolds, T.G.-ELT, VA-ELT, $\varepsilon^{2}$-PLT and $\varepsilon^{4}$-PLT solutions are compared against the Stokes solution (Appendix B). The accuracy of the pressure and velocity solutions is characterized using the relative percent error in $L_{2}$ norm. For an appropriate comparison of solutions from the various models, the flux $\mathcal{Q}$ is kept constant and the average pressure drops $\Delta P$ are presumed to vary.
Note, for the T.G.-ELT method, the prescribed flux $\mathcal{Q}$ may not be conserved throughout the domain – subject to the magnitude of surface variation – because incompressibility is not strictly enforced.
@@ 4.1 The logistic step
where $H_{\text{in}}\geq H_{\text{out}}>0$ and $\lambda>0$. The logistic step allows us to vary the magnitude of surface gradient while keeping a fixed length scale ratio $\varepsilon=L_{y}/L_{x}$ and step expansion ratio $H_{\text{in}}/H_{\text{out}}$. We consider solutions for $2\leq\lambda\leq 32$, and with fixed $H_{\text{in}}=2$, $H_{\text{out}}=1$, $L=16$, $\nu=1$, $\mathcal{U}=0$ and $\mathcal{Q}=1$. Then the length scale ratio is $\varepsilon\simeq 1/8$, and the maximum magnitude of surface gradient is $\big|\frac{dh}{dx}\big|=\lambda(H_{\text{in}}-H_{\text{out}})/4=\lambda/4$.
@@
Figure 2: Pressure and velocity solutions for the logistic step with $\lambda=8$. The ELT and $\varepsilon^{k}$-PLT solutions overestimate $\frac{\partial p}{\partial y}$ in the vicinity of the surface variation, leading to flow recirculation which is not observed in the Stokes solution at this moderate $\lambda$. The velocity for T.G.-ELT is not shown as it is equivalent to the Reynolds velocity.
The errors in velocity and pressure compared with the Stokes solution are shown in Fig. 3. For the wide range of $\lambda<32$, the VA-ELT solution of pressure is significantly more accurate than the Reynolds and T.G.-ELT solutions, and the T.G.-ELT solution of pressure yields a marginal improvement over the Reynolds pressure. For $2<\lambda<8$, the $\varepsilon^{2}$-PLT solution is best at approximating the pressure, and the VA-ELT solution is best at approximating the velocity. The $\varepsilon^{2}$-PLT solution performs better than the $\varepsilon^{4}$-PLT solution for all $\lambda>2$. For $\lambda\geq 32$, the Reynolds solution has the smallest error in velocity and pressure; the Reynolds solution has the most stable error across the range of $\lambda$. Overall, both the $\varepsilon^{2}$-PLT and VA-ELT solutions perform well for the logistic step up to moderate surface variation. We note however, that all the lubrication models discussed here break down in the limit $\lambda\to\infty$ of a discontinuous step, see Appendix D.
@@ 4.2 The triangular slider
(31) $h(x)=\begin{cases}H_{\text{in}}&x\leq L_{\text{in}}\\ H_{v}-\frac{H_{v}-H_{\text{in}}}{L_{a}}\big(L_{\text{in}}+L_{a}-x\big)&L_{\text{in}}\leq x\leq L_{\text{in}}+L_{a}\\ H_{\text{out}}-\frac{H_{\text{out}}-H_{v}}{L_{b}}\big(L-L_{\text{out}}-x\big)&L_{\text{in}}+L_{a}\leq x\leq L-L_{\text{out}}\\ H_{\text{out}}&L-L_{\text{out}}\leq x\leq L.\end{cases}$
We consider varying central vertex heights $\frac{1}{16}\leq H_{v}\leq 2$, and with fixed $L_{a}=1.25$, $L_{b}=0.75$, $L_{\text{in}}=L_{\text{out}}=7$, $H_{\text{in}}=H_{\text{out}}=1$, $\nu=1$, $\mathcal{U}=0$ and $\mathcal{Q}=1$. The maximum magnitude of surface gradient is $|\frac{dh}{dx}|=|1-H_{v}|/0.75$, and the length scale ratio is $\varepsilon\simeq\max(1,H_{v})/16$. This configuration is similarly considered in [tavakol_extended_2017] for the case of negative texturing $(H_{v}<1)$ with the $\varepsilon^{k}$-PLT solutions.
@@
The pressure and velocity solutions to the triangular slider with $H_{v}=0.5$ (following [tavakol_extended_2017]) and $H_{v}=2$ are shown in Figs. 5 and 6 respectively. This class of examples demonstrates the limitations of the lubrication models in the case of a discontinuous surface gradient. At the discontinuities in $\frac{dh}{dx}$, the pressure gradient $\frac{d^{2}p_{0}}{dx^{2}}$ and velocity $v_{0}$ of the Reynolds solution are not necessarily continuous. Consequently, the pressure and velocity solutions from ELT and $\varepsilon^{k}$-PLT may not be continuous at the discontinuities of $\frac{dh}{dx}$. Note that $u_{0}$ and $\frac{dp_{0}}{dx}$ are continuous for the triangular slider because $h(x)$ is continuous; owing to the lubrication assumptions, $v_{0}$ is typically small enough in magnitude compared to $u_{0}$ that the discontinuity in $v_{0}$ is not significant. However, for larger $H_{v}$, the discontinuity in $v_{0}$ is more pronounced and leads to a noticeable loss of smoothness in the ELT and $\varepsilon^{k}$-PLT solutions.
(a) Pressure contours $p(x,y)$
(b) Velocity streamlines $(u,v)$
@@
Figure 5: Pressure and velocity solutions for the triangular slider with $H_{v}=0.5$. The Reynolds, ELT and PLT models overestimate the maximum velocity. Discontinuities in velocity and pressure of the ELT and $\varepsilon^{k}$-PLT models are less pronounced at smaller $H_{v}$.
In the case of a sufficiently small interior angle at the triangle apex, the Stokes solution is known to exhibit a sequence of eddies receding into the corner [moffatt_viscous_1964, biswas_hoc_2017]. For $H_{v}=2$ as in Fig. 6, the first eddy is clearly observed in the Stokes solution. For the ELT and $\varepsilon^{k}$-PLT solutions, flow recirculation may occur if $H_{v}$ and $\frac{dh}{dx}$ are sufficiently large. However, because the velocities of these models are not continuous at the discontinuities of $\frac{dh}{dx}$, corner eddies do not occur as in the Stokes solution. In fact, as seen in Fig. 6 with the $\varepsilon^{k}$-PLT solutions, the velocity in the steeper half of the triangle texture exhibits flow recirculation at smaller $H_{v}$ than the shallower half. That is, if $H_{v}$ is large enough and $\frac{dh}{dx}$ is discontinuous in magnitude, the Reynolds, ELT and $\varepsilon^{k}$-PLT solutions may exhibit spontaneous flow reversal at the discontinuities of $\frac{dh}{dx}$.
(a) Pressure contours $p(x,y)$
(b) Velocity streamlines $(u,v)$
@@
Figure 6: Pressure and velocity solutions for the triangular slider with $H_{v}=2$. The velocity and pressure solutions of ELT and $\varepsilon^{k}$-PLT are discontinuous at the discontinuities of $\frac{dh}{dx}$. The sequence of corner eddies characteristic of the Stokes solution is not observed in any of the lubrication models.
@@
The errors in pressure and velocity compared with the Stokes solution for the triangular slider are shown in Fig. 7. The error in velocity is very similar for all of the models. For pressure, the VA-ELT solution is a consistent improvement compared to the Reynolds and T.G.-ELT solutions, and the $\varepsilon^{k}$-PLT solutions perform better still. For negative texturing ($H_{v}<1$), the $\varepsilon^{k}$-PLT solutions of pressure have significantly smaller error relative to both the VA-ELT and Reynolds solutions; for positive texturing ($H_{v}>1$), the $\varepsilon^{k}$-PLT and VA-ELT solutions are more similar. The error in both pressure and velocity decreases as the difference between the minimum and maximum height decreases. In particular, for the case of negative texturing, the increasing magnitude of surface gradient as the minimum height decreases corresponds to an increase in error for pressure and velocity; this emphasizes the importance of considering the magnitude of surface gradients as well as the length scale ratio when selecting a fluid model.
@@
Figure 7: Relative percent error in velocity $(u,v)$ and pressure $p(x,y)$ compared with the Stokes solution for the triangular slider with varying apex height $H_{v}$. Decreased difference between maximum and minimum height corresponds to decreased error in pressure and velocity.